OpenStudy (deadshot):

Solve for x: 4 over x plus 4 over quantity x squared minus 9 equals 3 over quantity x minus 3?

OpenStudy (anonymous):

4/x + 4/(x^2) - 9 = 3/(x - 3)? Did I do this right?

OpenStudy (deadshot):

\[\frac{ 4 }{ x }+\frac{ 4 }{ x^2-9 }=\frac{ 3 }{ x-3 }\]

OpenStudy (anonymous):

how else can you write \[x^2-9\]

OpenStudy (anonymous):

does the format ring bell?

OpenStudy (deadshot):

(x - 3i)(x + 3i), where i = √(-1), right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

where did i come from

OpenStudy (anonymous):

\[i= \sqrt -1\]

OpenStudy (anonymous):

it is just (x+3)(x-3)

OpenStudy (deadshot):

Oh

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now base is not equal do you know how to add numbers with unequal bases?

OpenStudy (deadshot):

no

OpenStudy (anonymous):

\[\frac{ a }{ b }+ \frac{ c }{ d } = \frac{ a*d+b*c }{ b*d }\]

OpenStudy (anonymous):

put the values tell me what you got

OpenStudy (deadshot):

\[\frac{ ((4x+12)(4x-12))+4x }{ (x^2+3x)(x^2-3x) }\]

OpenStudy (anonymous):

~_~ why did you multiply 4 to both the brackets

OpenStudy (deadshot):

I shouldn't have done that?

OpenStudy (anonymous):

no

OpenStudy (deadshot):

So only multiply 4 to the first bracket?

OpenStudy (anonymous):

yeah do this multiply by x^2-9

OpenStudy (anonymous):

keep the base value same

OpenStudy (anonymous):

which is x(x+3)(x-3)

OpenStudy (deadshot):

so, do I multiply 4 and (x^2-9)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i will save your work wait

OpenStudy (deadshot):

Okay

OpenStudy (anonymous):

\[\frac{ 4x^2+4x-12 }{ x(x+3)(x-3) } =\frac{ 3 }{ x-3 }\]

OpenStudy (anonymous):

now take 4 common you will get 4(x^2+x-12)

OpenStudy (anonymous):

factorize the term in bracket then you will get 4(x-1)(x+3)

OpenStudy (deadshot):

And now I factor it to \[4(x-3)(x+4)\]

OpenStudy (anonymous):

\[\frac{ 4(x-1)(x+3) }{ x(x+3)(x-3) } =\frac{ x }{ (x-3) }\]

OpenStudy (anonymous):

hold on

OpenStudy (deadshot):

So then, it's \[\frac{ 4(x-3)(x+4) }{ x(x+3)(x-3) }=\frac{ x }{ (x-3) }\]?

OpenStudy (anonymous):

4*9 = 36 ~_~ sorry i got the equation wrong

OpenStudy (deadshot):

okay, so what would the correct equation be?

OpenStudy (anonymous):

\[4x^2+4x-36\]

OpenStudy (deadshot):

okay

OpenStudy (anonymous):

so take out 4 you will get x^2+x-9 now solve

OpenStudy (deadshot):

factorized, it is (x-3)(x+4)

OpenStudy (anonymous):

You don't want to expand anything. Keeping things factored makes it easy to see roots.

OpenStudy (anonymous):

(x-3)(x+4) ? = x^2+x+12

OpenStudy (anonymous):

-12 **

OpenStudy (anonymous):

Nevermind, you will have to expand at some point.

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

there are no real roots for x^2+x-9

OpenStudy (deadshot):

So x²+x-9 is not factorable under the reals, however it shouldbr factorable under the complex numbers. If i'm correct, then (x^2+x-9)=(x - 3i)(x + 3i), where i = √(-1) right?

OpenStudy (deadshot):

*should be

OpenStudy (anonymous):

LOL dont go for complex numbers this is not related to complex numbers

OpenStudy (deadshot):

okay

OpenStudy (deadshot):

So complex numbers would just complicate it more

OpenStudy (anonymous):

here is the solved 1 when you cancel equal factors x-3 you will get \[4x^2+4x-36=3x(x+3)\]

OpenStudy (anonymous):

last equation will be \[x^2-5x-36=0\]

OpenStudy (anonymous):

which is \[(x-9)(x+4)=0\]

OpenStudy (anonymous):

so your answer is x= 9 or x= -4

OpenStudy (deadshot):

Okay, Thanks!

OpenStudy (anonymous):

you are welcome @wio thanks i would hv got lost factorizing this :D